OpenStudy (anonymous):

A roller coaster has a mass of 275 kg. It sits at the top of a hill with height 85 m. If it drops from this hill, how fast is it going when it reaches the bottom? (Assume there is no air resistance or friction.) A. 18.0 m/s B. 29.9 m/s C. 73.4 m/s D. 40.8 m/s Just need to know how to work problems that are similar.

OpenStudy (rane):

when the roller coaster is on top of the hill, it has GRAVITATIONAL POTENTIAL ENERGY. bt when it reaches the bottom that GPE coverts to KINETIC ENERGY

OpenStudy (rane):

so u have to find the GPE and wht ever is your answer..that would also equal to KE PE=KE

OpenStudy (rane):

EP= mgh

OpenStudy (rane):

so, its 275*9.8*85

OpenStudy (anonymous):

WEll you can here conserve the total mechanicl energy so at the top,total mech. E=mgh and at the bottom its =1/2m[v ^{2}] also you can use newton's third law i.e 2gh=(v-u)(v+u),,where initial vel u=0 at the top and v=final vel. at bottom

OpenStudy (rane):

now just calculate the answer

OpenStudy (anonymous):

Thanks so much(: I don't understand physics and I just need to finish this class online to graduate in 2 weeks.

OpenStudy (anonymous):

@RANE i think the ans is(2*9.8* 85)^1/2

OpenStudy (rane):

PE= mgh

OpenStudy (rane):

where did u get 1/2 from?

OpenStudy (anonymous):

well 1/2mv^2=mgh

OpenStudy (rane):

yes bt right now u r looking for PE

OpenStudy (rane):

when the roller coater will move down it will transfer its PE to KE because its no more above the ground level

OpenStudy (anonymous):

With the (1/2) it comes to 833. That isn't right

OpenStudy (rane):

exactly, u have to use PE @DaniRae231

OpenStudy (anonymous):

well its not half its under root.. ans is 41.smthng

OpenStudy (anonymous):

40.8 m/s

OpenStudy (rane):

whts the answer given to u ? @DaniRae231

OpenStudy (anonymous):

Its an online class. I answer it and it tells me if I got it right or not. It's APEX.

OpenStudy (anonymous):

@DaniRae231 yes i too got that

OpenStudy (rane):

u can not use EK formula when u r supposed to find PE becoz its above the ground level...and also its not moving

OpenStudy (rane):

if an object is not moving then how can u use kinetic energy

OpenStudy (anonymous):

well we have to find the final velocity....so we conserve the mech. energy at the top its only mgh( vel =0) and at the bottom its 1/2mv^2(h =0)....so mgh =1/2mv^2

OpenStudy (anonymous):

@RANE ,i didnt get u

OpenStudy (rane):

ok... when do u calculate PE

OpenStudy (rane):

i mean when its moving or its above the ground level

OpenStudy (anonymous):

well potential energy ,when it is above the ground

OpenStudy (rane):

exactly, now do u know that energy is never lost right?

OpenStudy (anonymous):

perfectly

OpenStudy (rane):

wait wht grade r u in?

OpenStudy (anonymous):

m in 12th stndrd..why?

OpenStudy (rane):

no, just asking..anyway if roller coaster is above the ground at rest it has PE bt as it starts to come doen that PE starts to covert in to kinetic energy

OpenStudy (anonymous):

yes or we can say pot energy decreases and kinetic energy starts to increase...

OpenStudy (rane):

yes

OpenStudy (anonymous):

Well I already have it, so it doesn't matter to me...

OpenStudy (anonymous):

and in which grade u are @RANE ?

OpenStudy (rane):

so since there is no lost in energy then PE= KE if u calculate the potential energy (since it's mgh and m and h is given to u) it will equal to kinetic enrgy at the bottom

OpenStudy (anonymous):

yes i also did the same..you are very correct

OpenStudy (rane):

umm. so it means that u understood y we cant use 1/2 in calculation

OpenStudy (anonymous):

Alright, this is just a Junior science Physics class. I just failed it my junior year, so they have me doing it online.

OpenStudy (rane):

this is just the basic of energy conservation in phys...

OpenStudy (anonymous):

no...i jus actually said that potential energy at the top(mgh)=kinetic energy at the bottom(1/2m(v)^2)...so 1/2 will come for sure

OpenStudy (rane):

no!!!

OpenStudy (rane):

omg

OpenStudy (anonymous):

why..i want to know!

OpenStudy (rane):

now i will just recommend u to use google or ask your teacher

OpenStudy (anonymous):

well tell me coz i think m correct

OpenStudy (rane):

LOL no u r not correct

OpenStudy (rane):

just go nd study conservation of energy

OpenStudy (anonymous):

see at the top total mech energy=mgh + 1/2m(0)^2 =mgh + 0 nw at the bottom total mech energy=mg(0) +1/2mv^2 hence they will b conserved...sry i think u got it totally wrong @RANE..n the ans is also there

OpenStudy (rane):

noooooo........ i've dont these for one hole term this yr which is only 4 wks ago when i did my final examination on it so u can not tell me that i got my answers rong in my exam and top of that the exam markers also marked them wrong

OpenStudy (anonymous):

only one way to sort it out...how did u do it?